# Resisting Contextual Interference in RAG via Parametric-Knowledge Reinforcement

### v2

@@ 1 Introduction
When contextual knowledge appears reliable but is actually incorrect, LLMs should ignore it and decide whether to fall back on parametric knowledge. This behavior is rare but crucial, as current LLMs tend to rely more on context when faced with conflicting or misleading information. Reinforcement learning (RL) can assist by adjusting the probability rankings of outputs, encouraging the LLM to explore these rare but critical decisions, such as when to fall back on parametric knowledge (Stiennon et al., 2020; Ramamurthy et al., 2023). Therefore, we believe RL has the potential to help LLMs learn how to effectively balance and utilize both parametric and contextual knowledge. Recently, models like OpenAI-o1 (Jaech et al., 2024) and DeepSeek-R1 (Guo et al., 2025) have employed RL techniques such as PPO (Schulman et al., 2017) and GRPO (Shao et al., 2024) to enhance LLMs’ logical reasoning and problem-solving abilities through experience and feedback. However, in RAG systems, these methods are limited by their sampling space and objectives, which leads to a primary focus on context-aware reasoning while overlooking the importance of incorporating parametric knowledge.
@@ 4.1.1 Reasoning Instruction and Baselines
We evaluate with Qwen2.5-3B-Instruct, Qwen2.5-7B-Instruct, and Llama3.1-8B-Instruct. Baselines include query-only prompting, RAG prompting, Astute-RAG (a prompting method for imperfect retrieval (Wang et al., 2025a)), and GRPO with RAG. We evaluate using exact match (EM) (Jin et al., 2025). To ensure fairness, all methods use the same total training data and are trained for one epoch. For GRPO and Knowledgeable-R1, we use a global batch size of 128, a rollout batch size of 32, a rollout temperature of 1, and a learning rate of $1\times 10^{-6}$. All experiments run on 8 H100 GPUs, using the same system prompt for both training and inference, where the inference input combines the query and context. For further experiment details, see Appendix B, C, and D.
@@ 4.2.1 Performance across Contextual Scenarios
When the context has internal contradictions, Knowledgeable-R1 performs best among both model families: 63.77% compared to 63.38% for GRPO w/ RAG on qwen2.5-7B, and 76.58% compared to 73.67% on llama3.1-8B. Though the improvements are small, the consistent gains across models show stable handling of conflicting information. Resolving contradictions within the context seems difficult for all methods, indicating an area for future improvement.
@@
In the ExplainPE benchmark, where the context is completely unrelated, we see the typical "RAG-hurt" effect: query-only (64.45%) performs better than RAG prompting (62.21%) on qwen2.5-7B. Knowledgeable-R1 achieves 67.57%, actively rejecting irrelevant context and outperforming both methods. This shows that the method can recognize when context isn’t helpful for the query.
@@
When valid evidence is mixed with distractors, Knowledgeable-R1 performs strongest or nearly strongest. On qwen2.5-7B, it achieves 31.45% on HotpotQA, 37.52% on 2WikiMultiHopQA, and 12.04% on MuSiQue, with significant gains over RAG prompting. Notably, the strong performance on 2WikiMultiHopQA and MuSiQue is achieved without fine-tuning on these datasets, showing that the method can generalize to new evidence types and distractor patterns beyond the training data (HotpotQA). This ability to generalize across datasets is especially useful for real-world applications, where the evidence may differ from the training data.
@@ Appendix A Prompt designed in Knowledgeable-R1 and baselines
Fine-tuning-based Methods: We evaluate Knowledgeable-R1 against GRPO (Guo et al., 2025) with RAG prompting (GRPO $w/$ RAG).

### v3

@@ 1 Introduction
When contextual knowledge appears reliable but is actually incorrect, LLMs should ignore it and decide whether to fall back on parametric knowledge. This behavior is rare but crucial, as current LLMs tend to rely more on context when faced with conflicting or misleading information. Reinforcement learning (RL) can assist by adjusting the probability rankings of outputs, encouraging the LLM to explore these rare but critical decisions, such as when to fall back on parametric knowledge (Stiennon et al., 2020; Ramamurthy et al., 2023). Therefore, we believe RL has the potential to help LLMs learn how to effectively balance and utilize both parametric and contextual knowledge. Recently, models like OpenAI-o1 (Jaech et al., 2024) and DeepSeek-R1 (Guo et al., 2025b) have employed RL techniques such as PPO (Schulman et al., 2017) and GRPO (Shao et al., 2024) to enhance LLMs’ logical reasoning and problem-solving abilities through experience and feedback. However, in RAG systems, these methods are limited by their sampling space and objectives, which leads to a primary focus on context-aware reasoning while overlooking the importance of incorporating parametric knowledge.
@@ 4.1.1 Reasoning Instruction and Baselines
We evaluate with Qwen2.5-3B-Instruct, Qwen2.5-7B-Instruct, Qwen2.5-14B-Instruct and Llama3.1-8B-Instruct. Baselines include query-only prompting, RAG prompting, Astute-RAG (a prompting method for imperfect retrieval (Wang et al., 2025a)), CK-PLUG(Bi et al., 2025c) and GRPO(Guo et al., 2025a) with RAG. We evaluate using exact match (EM) (Jin et al., 2025). To ensure fairness, all methods use the same total training data and are trained for one epoch. For GRPO and Knowledgeable-R1, we use a global batch size of 128, a rollout batch size of 32, a rollout temperature of 1, and a learning rate of $1\times 10^{-6}$. All experiments run on 8 H100 GPUs, using the same system prompt for both training and inference, where the inference input combines the query and context. For further experiment details, see Appendix B, C, and D.
@@ 4.2.1 Performance across Contextual Scenarios
When the context has internal contradictions, Knowledgeable-R1 performs best among both model families: 63.77% compared to 63.38% for GRPO w/ RAG on Qwen2.5-7B-Instruct, and 76.58% compared to 73.67% on llama3.1-8B. Though the improvements are small, the consistent gains across models show stable handling of conflicting information. Resolving contradictions within the context seems difficult for all methods, indicating an area for future improvement.
@@
In the ExplainPE benchmark, where the context is completely unrelated, we see the typical "RAG-hurt" effect: query-only (64.45%) performs better than RAG prompting (62.21%) on Qwen2.5-7B-Instruct. Knowledgeable-R1 achieves 67.57%, actively rejecting irrelevant context and outperforming both methods. This shows that the method can recognize when context isn’t helpful for the query.
@@
When valid evidence is mixed with distractors, Knowledgeable-R1 performs strongest or nearly strongest. On Qwen2.5-7B-Instruct, it achieves 31.45% on HotpotQA, 37.52% on 2WikiMultiHopQA, and 12.04% on MuSiQue, with significant gains over RAG prompting. Notably, the strong performance on 2WikiMultiHopQA and MuSiQue is achieved without fine-tuning on these datasets, showing that the method can generalize to new evidence types and distractor patterns beyond the training data (HotpotQA). This ability to generalize across datasets is especially useful for real-world applications, where the evidence may differ from the training data.
@@ Appendix A Prompt designed in Knowledgeable-R1 and baselines
Fine-tuning-based Methods: We evaluate Knowledgeable-R1 against GRPO (Guo et al., 2025b) with RAG prompting (GRPO $w/$ RAG).